# A geometric approach toconjugation-invariant random permutations

### v3

@@ 7.4 Cycle types with few fixed points
Since the family (.$\left(\big{.}(u,v)\mapsto g_{i}(u)g_{j}(v)\right)_{1\leq i,j\leq r}$ is linearly independent in $L^{2}([0,1]^{2})$, the rank of $(\chi_{\pi})_{\pi\in\mathfrak{S}_{r}}$ is equal to the rank of the family $\left(\alpha A^{\pi}+\beta A^{(\pi^{-1})}\right)_{\pi\in\mathfrak{S}_{r}}$ in $\mathcal{M}_{r}(\mathbb{R})$. As explained in [JNZ15] (see the proof of Theorem 2.5 therein), the family $(A^{\pi})_{\pi\in\mathfrak{S}_{r}}$ spans the space of matrices with rows summing to 0 and columns summing to 0, and thus has rank $(r-1)^{2}$.
@@
Now suppose $p_{2}=1$ i.e. $\alpha=\beta$. It is easily checked that the family $\left(A^{\pi}+A^{(\pi^{-1})}\right)_{\pi\in\mathfrak{S}_{r}}$ spans the space of symmetric matrices with rows summing to 0 and columns summing to 0. Such a matrix $A$ is exactly characterized by its coefficients $A_{i,j}$ with $i\leq j<r$, hence this space has dimension $r(r-1)/2$. This concludes the proof. ∎
Acknowledgements
The author is very grateful to his advisor Valentin Féray for insightful discussions and valuable suggestions, and to Mohamed Slim Kammoun for helpful comments.

### v4

@@ 7.4 Cycle types with few fixed points
Since the family $\left(\big.(u,v)\mapsto g_{i}(u)g_{j}(v)\right)_{1\leq i,j\leq r}$ is linearly independent in $L^{2}([0,1]^{2})$, the rank of $(\chi_{\pi})_{\pi\in\mathfrak{S}_{r}}$ is equal to the rank of the family $\left(\alpha A^{\pi}+\beta A^{(\pi^{-1})}\right)_{\pi\in\mathfrak{S}_{r}}$ in $\mathcal{M}_{r}(\mathbb{R})$. As explained in [JNZ15] (see the proof of Theorem 2.5 therein), the family $(A^{\pi})_{\pi\in\mathfrak{S}_{r}}$ spans the space of matrices with rows summing to 0 and columns summing to 0, and thus has rank $(r-1)^{2}$.
@@
Now suppose $p_{2}=1$ i.e. $\alpha=\beta$. It is easily checked that the family $\left(A^{\pi}+A^{(\pi^{-1})}\right)_{\pi\in\mathfrak{S}_{r}}$ spans the space of symmetric matrices with rows summing to 0 and columns summing to 0. Such a matrix $A$ is exactly characterized by its coefficients $A_{i,j}$ with $i\leq j<r$, hence this space has dimension r(r1)/2r(r-1)/2. This concludes the proof. ∎